OpenStudy (anonymous):

factor completely: see below:

OpenStudy (anonymous):

\[125m ^{3}-8x ^{3}\] = \[(5m-2x)(25m ^{2}+10mx+4x ^{2}\] can someone please explain how to get the +10mx???????????????

OpenStudy (anonymous):

I can't figure out how to get the middle number??

OpenStudy (zzr0ck3r):

\[a^3 – b^3 = (a – b)(a^2 + ab + b^2)\\you \space have\space a=5m, b=2m\\so\\125m^3-8x^3 = ((5m)^3)-(2m)^3)=(5m-2x)(25m^2+10mx+4x^2)\]

OpenStudy (anonymous):

how did you figure middle number? 10mx

OpenStudy (anonymous):

ok got it